OpenStudy (anonymous):

The equation of the tangent line to the curve x^2+y^2=169 at the point (5,-12) is

myininaya (myininaya):

first to find the slope of the tangent line you need y'

myininaya (myininaya):

2x+2yy'=0 x+yy'=0 y'=-x/y now to find slope just plug 5 in for x and -12 for y

myininaya (myininaya):

y=mx+b y=-5x/12+b we know a point on this line so we can find b -12=-5(5)/12+b solve for b and then we will have equation of the line

myininaya (myininaya):

b=-12+25/12=-119/12 y=-5x/12-119/12

OpenStudy (anonymous):

Thanks! :)

myininaya (myininaya):

your welcome